William bought shares 20.3. and sold them 6.6. in the same year. The purchase price was 473 € and the selling price 639 €. What is the annual interest rate for the investment after the 30 % tax is paid from the sales profit?


1
Expert's answer
2022-02-08T13:56:46-0500

Gross profit "=639-473=166\u20ac"

Net profit "=\frac{70}{100}×166=116.2"


Let r "=" annual interest rate


Number of days from 20th March to 6th June "=78" days


"\frac{r}{100}×\frac{78}{365}×473=116.2"


"r=114.96\%"